# FP-Inconsistent

### v1

@@ 7.3 Accuracy improvement through inconsistency analysis
Our results on the requests received on our honey site show that using a filter list to counter commonly found inconsistencies is an effective method to detect and block evasive bots. Filter lists are commonplace in anti-tracking community, where they provide a good trade off between performance and accuracy in detecting advertising and tracking services. Currently, no such alternative exist to detect bots that show inconsistent fingerprints. Our methodology, described in Algorithm 1, is a first step towards creating such filter list to enhance online bot detection.
8 Discussion
8.1 Improving FP-Inconsistent
8.1.1 Capturing more attributes
@@
Even if we assume that bots have the ability to alter all browser attributes, inconsistencies provide a promising avenue for detection as long as there exist at least one pair of attributes that cannot exist in the real world. Accordingly, increasing the number of captured attributes introduces more avenues for inconsistencies among those attributes which can be leveraged for detection. In this paper, we confined FP-Inconsistent to only look for inconsistencies among HTTP headers and the attributes captured by FingerprintJS. Incorporating other browser attributes such as those captured by CreepJS can further improve FP-Inconsistent’s detection accuracy. Incorporating behavioral characteristics of users can also bolster bot detection. However, capturing such characteristics to improve bot detection comes at the cost of privacy since these attributes enable tracking when captured from real users. Privacy researchers seek to limit the amount of information exfiltrated by browsers to protect privacy.
8.1.2 Capturing attributes from real users
In this paper, we used the requests sent by bots on our honey site to devise rules for inconsistencies. Thus, these rules can be further refined by incorporating information about browser attributes captured from requests from real users. However, our honey site architecture can only guarantee reliable ground-truth that requests originate from bots and cannot be extended to guarantee requests originating from real users. Ensuring reliable ground-truth that requests originate from real users is difficult without a public deployment that naturally attracts diverse users to the site. Such a deployment still risks attracting bots thereby making it harder to guarantee ground-truth.
8.1.3 Capturing unmodifiable attributes
@@
Researchers have proposed side-channels to capture fundamental physical device characteristics to uniquely identify devices even among those that have identical hardware and software configurations [44, 45, 46, 47]. Such techniques can significantly empower temporal inconsistencies to detect bots. In our inconsistency analysis, we used Cookies to identify requests that originated from the same device. Bots will be able to overcome our temporal inconsistencies by merely deleting their cookies. Bots would not be able to drop unique identifiers that originate from the physical properties of hardware that cannot be modified. Such persistent identifiers, however, pose threats to privacy.
8.2 Deployment of filter list rules
As described in section 7.3, FP-Inconsistent can be used to generate a filter list of inconsistencies to improve bot detection. Use of filter lists to protect against trackers is a widely used approach in anti-tracking community [48, 49, 50, 51, 52, 53] These filter lists are used on the client side via an extension that matches outgoing requests and other resources with the filter lists rules to block tracking requests and other resources from executing. The filter list produced by FP-Inconsistent can be used in a similar fashion However, because the filter rules proposed by FP-Inconsistent are meant to aid in detection of evasive bots, we envision the inclusion of these rules in client-side scripts loaded by anti-bot services such as DataDome and BotD. As per our analysis in section 4.2, these services use client side scripts to collect fingerprints to aid in detection, which are then relayed to their servers for a final decision. Inclusion of filter list rules generated by FP-Inconsistent on the the client side will help these services improve detection of evasive bots.
8.3 Limitations
8.3.1 False positives from privacy-enhancing technologies
Privacy enhancing technologies [54, 55, 56] alter browser attributes to protect users from being tracked online. Thus, requests sent by users employing such technologies may present inconsistent attributes resulting in them being flagged as bots by our inconsistency analysis.
8.3.2 Evasive bots with consistent attributes
@@
Our results show that FP-Inconsistent’s inconsistency rules can improve the detection of evasive bots that are currently evading detection. Evasive bots will be able to overcome FP-Inconsistent if they evolve to ensure that they can alter attributes without introducing any inconsistencies. Incorporating unmodifiable attributes can help provide a fundamental solution to detect bots, but doing so comes at the cost of privacy.
8.4 Need for privacy-preserving bot detection
@@
Our discussion in this section reveal a forceful trade-off between improving bot detection and being respectful of user privacy. Orthogonal to the scope of this paper, future research focusing on privacy-preserving bot detection such as identifying intent behind trackers to not block those interested in bot detection or an in-browser detection mechanism can bridge the gap to potentially address concerns of privacy protection as well as bot detection.
9 Conclusion
@@
We find evidence that bots alter browser fingerprint attributes to evade detection. However, we find evidence that such evasive bots end up introducing inconsistencies among the fingerprint attributes that can be used for more reliable bot detection. We propose FP-Inconsistent, a data-driven, semi-automatic approach to discover inconsistencies in browser fingerprint attributes for detecting evasive bots in the wild that are able to evade detection by anti-bot services. As the arms race between evasive bots and anti-bot services evolves, it remains to be seen whether bots can alter their browser fingerprint attributes while avoiding inconsistency. We believe that it would be challenging for bots to do so because a browser fingerprint is a high dimensional feature set with numerous – often subtle – correlations between attributes that are difficult to anticipate and account for when altering fingerprints. Put simply, it is challenging to tell a complex lie while keeping the story always straight. While FP-Inconsistent rule generation approach may need to be evolved to generate rules for other types of consistencies for future generation of bots, we believe the basic principle will stand over time.
Appendix A Comparison of APIs used by BotD and DataDome
@@
Table V shows the different APIs accessed by BotD and DataDome scripts on our honey site.
TABLE V: Comparison of browser APIs read by DataDome and BotD
@@
Appendix B Inconsistencies Identified
Table VI lists some examples of the inconsistencies that we identified for each feature group in Table VII.
TABLE VI: Inconsistencies Identified
Feature Group Features Examples
@@ Appendix B Inconsistencies Identified
TABLE VII: Feature Categories
Category Features
@@
Appendix C Feature Categories for Inconsistency analysis
Table VII list different categories of features used for inconsistency analysis.
Figure 14: Pie chart showing the proportion of requests originating from IP addresses that are not present in block lists, that are present in block lists but evading and detected by DataDome.
Figure 15: Pie chart showing the proportion of requests originating from IP addresses that are not present in block lists, that are present in block lists but evading and detected by BotD.
Appendix D IP block lists
Since ASNs are coarser than IP addresses, we directly check if evasive bots send requests from blocked IP addresses. Concretely, we queried the IP addresses of the requests received on our honey site against MaxMind’s minFraud API  [38]. Figure shows that 15.86% of requests originated from IP addresses that were blocked by MaxMind. Among the requests that originated from blocked IP addresses, 48.1% of requests were not detected by DataDome and 68.85% requests were not detected by BotD. Requests originating from blocked IP addresses being able to evade detection indicates that evasive bots don’t merely rely on IP addresses for evasion.
Appendix E Combination of browser attributes to evade DataDome

### v2

@@ 7.3. Improved bot detection
Our results on the requests received on our honey site show that using a filter list to counter commonly found inconsistencies is an effective method to detect and block evasive bots. Filter lists are commonplace in the anti-tracking community, where they provide a good trade-off between performance and accuracy in detecting advertising and tracking services. Currently, no such alternative exists to detect bots that show inconsistent fingerprints. Our methodology is a first step towards creating such filter lists to enhance online bot detection.
7.4. Real user traffic
We also evaluate FP-Inconsistent’s filter rules against traffic from real users to ensure that our improvements in bot detection do not incorrectly detect real users as bots. Concretely, we shared a version of our honey site that contained a unique URL with students at our university. Since we only shared this URL with bonafide students, we have high confidence that requests from real users were recorded at this URL. We did not collect any Personally Identifiable Information (PII) from these users and discuss the ethics of collecting this data in Appendix A.
We report a true negative rate of 96.84% on the 2,206 requests received at this URL. The small number of false positives were likely due to students experimenting with User-Agent spoofers, as these cases triggered spatial inconsistencies involving User-Agents. We could not conduct large-scale evaluation on real user traffic in the wild since it would be challenging to ensure ground-truth. Regardless, our evaluation shows low false positive rates, which can be further mitigated using CAPTCHAs if needed (Section 8.1).
7.5. Privacy-enhancing browsers
Privacy-enhancing browsers such as Brave (Brave, [n. d.]), Tor (Tor, [n. d.]), and Fingerprint Spoofer (Fingerprint Spoofer, [n. d.]) alter fingerprint attributes to protect user privacy against tracking (Iqbal et al., 2020; Ikram et al., 2016). In this section, we examine the attributes altered by such technologies and their impact on FP-Inconsistent.
We conducted an experiment where we sent requests to different versions of our honey site (each with a distinct URL) while employing five different privacy-enhancing browsers: Safari, Brave, Tor browsers as well as uBlock Origin and AdBlockPlus browser extensions on Google Chrome. We 300 requests from devices running macOS (M1 MacBook Pro), Linux (Intel Coffee Lake Desktop), iOS (iPad Pro), and Android (Google Pixel 7).
Brave Brave browser currently alters 6 different fingerprint attributes: audio, canvas, plugins, deviceMemory, hardwareConcurrency, and screenResolution. Our inconsistency rules do not incorporate the former three attributes and Brave’s alterations to the others were consistent with other attributes. For instance, Brave alters deviceMemory on desktops to plausible values (0.5, 1, 2, 4, and 8), which align with the amount of memory in typical desktops and remain consistent with other fingerprint attributes.
However, since Brave browser retains Cookies across requests, the requests triggered several temporal inconsistencies where multiple requests shared the same Cookie but had differing values for both hardwareConcurrency and deviceMemory. Such inconsistencies are rare in real-world scenarios, as they require users to enable Brave’s fingerprint protection while retaining Cookies. These rare false positives can be mitigated using CAPTCHAs, with the verification result stored in Cookies (Section 8.1).
Although FP-Inconsistent does not detect requests from Brave browser as bots, we argue that bot services cannot exploit Brave for evasion since they seek to alter attributes that are not supported by Brave. Concretely, we see that Brave only alters 2 attributes that are most commonly altered by evasive bots (Section 5.2) and does not alter other attributes that are of interest to bots such as those pertaining to their device type or geolocation (Section 6). If Brave were to alter more fingerprint attributes in the future, FP-Inconsistent could become prone to more false positives. However, even in such a hypothetical scenario, these false positives can be mitigated using CAPTCHAs (Section 8.1. Moreover, only a small set of users would encounter these CAPTCHAs if Brave’s market share continues to remain at 1% (TechReport, [n. d.]).
Tor FP-Inconsistent detected all requests from Tor browser as bots since they triggered spatial inconsistencies between the geolocation inferred from their IP address and the timezone attribute of their navigator object. While Tor results in false positives, we expect a small set of users to be affected since Tor likely has less than 1% market share (StatCounter, 2024). Furthermore, most websites currently block requests from Tor (Tor, [n. d.]), due to the difficulty in distinguishing Tor traffic from bots. To mitigate false positives, we can present users with CAPTCHAs rather than blocking their requests (Section 8.1). We report the detection accuracy of DataDome and BotD on Brave and Tor traffic in Appendix G.
Safari, uBlock Origin, and AdBlockPlus None of these requests were detected as bots. This is because these tools protect privacy by blocking tracking requests rather than altering fingerprint attributes. The two extensions cater to over 80 million users combined on the most widely used browser, Google Chrome (StatCounter, 2024; Webstore, 2024b, a). Safari has the second largest market share among web browsers (StatCounter, 2024). This shows that FP-Inconsistent can detect bots while having zero impact of all these users.
8. Discussion
8.1. Overcoming false positives
Our evaluation on requests from real users show that FP-Inconsistent incurs low, but non-zero false positive rates (Section 7.4). Our experiments with privacy-enhancing technologies also reveal certain scenarios that could lead to false positives (Section 7.5). In the context of this paper, false positives refer to requests from real users that were incorrectly detected as bots. Challenging users to solve CAPTCHAs rather than blocking them offers a promising solution to mitigate false positives (Amin Azad et al., 2020; Din et al., 2020). While effective, CAPTCHAs could potentially frustrate certain users (Nakatsuka et al., 2021; Searles et al., 2023). This frustration can be mitigated by storing the result of a CAPTCHA verification in a Cookie, thereby reducing the frequency at which users are asked to solve CAPTCHAs.
8.2. Improving FP-Inconsistent
@@
Inconsistencies provide a promising avenue for detection as long as there exist at least one pair of attributes that cannot exist in the real world. Accordingly, increasing the number of captured attributes introduces more opportunities for inconsistencies which can be leveraged for detection. In this paper, we confined FP-Inconsistent to only look for inconsistencies among HTTP headers and the attributes captured by FingerprintJS. Incorporating other attributes such as those from CreepJS (abrahamjuliot, [n. d.]) can further improve FP-Inconsistent.
@@
Researchers have proposed side-channels based on physical device characteristics to uniquely identify devices even among those with identical hardware and software configurations (Sanchez-Rola et al., 2018; Laor et al., 2022; Venugopalan et al., 2023; Schaller et al., 2017). Such techniques can significantly empower temporal inconsistencies to detect bots. With FP-Inconsistent, we used Cookies to identify requests that originated from the same device. Bots will be able to overcome our temporal inconsistencies by merely deleting their cookies. Bots would not be able to drop unique identifiers that originate from the physical properties of hardware that cannot be modified. However, capturing more attributes as well as capturing persistent identifiers pose threats to privacy.
8.3. Deployment of filter list rules
FP-Inconsistent generates filter lists of inconsistencies to improve bot detection (Section 7.3). The anti-tracking community (Siby et al., 2022; Iqbal et al., 2018; Munir et al., 2023; Inc., [n. d.]; AdguardTeam, [n. d.]; gorhill, [n. d.]) typically incorporates filter lists on the client side using browser extensions to block the execution of tracking requests and other resources. Similarly, we envision anti-bot services such as DataDome and BotD to include FP-Inconsistent’s filter lists as part of their client-side scripts improved for bot detection.
8.4. Limitations
@@
Our results show that FP-Inconsistent’s rules improve the detection of evasive bots. Evasive bots will be able to overcome FP-Inconsistent if they evolve to ensure that they can alter fingerprint attributes without introducing any inconsistencies. Incorporating unmodifiable attributes provides a robust solution to enhance FP-Inconsistent, but such attributes also pose threats to privacy.
8.5. Coexistence of Bot Detection and Privacy-Enhancing Technologies
Our evaluation (Section 7.5) shows that FP-Inconsistent can detect bots without impeding most commonly used privacy-enhancing technologies (except Tor). This observation is interesting because many assume that the goals of bot detection and online tracking are identical. They believe that enhancements to bot detection would also bolster online tracking and weaken privacy-enhancing tools.
While bot detection and tracking overlap, tracking is more complex as it seeks to uniquely identify each user. In contrast, bot detection solely seeks to determine if a particular request was generated by a bot. Accordingly, altering any fingerprint attribute enhances privacy by making it harder for trackers to link requests from the same user, even when other attributes remain unchanged. On the other hand, altering any fingerprint attribute does not necessarily help bots with evasion since other attributes can still reveal their presence. This distinction between bot detection and online tracking allows bot detection systems like FP-Inconsistent to coexist with privacy-enhancing technologies.
@@ 8.5. Coexistence of Bot Detection and Privacy-Enhancing Technologies
However, given the overlap, certain enhancements to bot detection such as incorporating more attributes or incorporating unmodifiable attributes can threaten user privacy. Future research focusing on privacy-preserving bot detection such as identifying the intent behind trackers to not block those indulging in bot detection or an in-browser detection mechanism can bridge the gap to potentially address concerns of privacy protection as well as bot detection.
9. Conclusion
@@
We find evidence that bots alter fingerprint attributes to evade detection. However, we find evidence that such evasive bots end up introducing inconsistencies among the fingerprint attributes that can be used for more reliable bot detection. We propose FP-Inconsistent, a data-driven, semi-automatic approach to discover inconsistencies in fingerprint attributes for detecting evasive bots in the wild that are able to evade detection by anti-bot services. As the arms race between evasive bots and anti-bot services evolves, it remains to be seen whether bots can alter their fingerprint attributes while avoiding inconsistency. We believe that it would be challenging for bots to do so because a browser fingerprint is a high dimensional feature set with numerous – often subtle – correlations between attributes that are difficult to anticipate and account for when altering fingerprints. Put simply, it is challenging to tell a complex lie while keeping the story always straight. While FP-Inconsistent rule generation approach may need to be evolved to generate rules for other types of consistencies for future generation of bots, we believe the basic principle will stand over time.
Appendix A Ethics
This study complies with ethical guidelines for research involving data collection and usage. To conduct the research, we paid a small amount to bot operators to generate requests directed solely to our honey site. To ensure data quality and realism, we prioritized bot services with high ratings and traffic advertised as realistic and organic. These requests were analyzed exclusively for research purposes, with the goal of improving bot detection.
The research process was reviewed and approved by our university, ensuring alignment with ethical principles outlined in both the Belmont Report and the Menlo Report. To determine whether Institutional Review Board (IRB) approval was necessary, we consulted official guidelines from the Human Subject Regulations Decision Charts (U.S. Department of Health and Human Services, 2018), specifically the section addressing activities covered by 45 CFR Part 46. Based on this evaluation, we determined that our research does not involve human subjects as defined by 45 CFR Part 46 and, as a result, qualifies for exemption from IRB oversight.
Furthermore, our study did not collect or store any Personally Identifiable Information (PII), nor did it involve the identification or tracking of individual users across different websites/contexts. Traffic data was analyzed in aggregate, and identifiable information, such as IP addresses, was hashed before storage.
All purchased traffic was directed exclusively towards our honey site, ensuring that no other sites or users were impacted. The primary purpose of this research was to advance the science of bot detection, and we refrained from monetizing the honey site or deriving any profit from the generated traffic.
Appendix B Comparison of APIs used by BotD and DataDome
@@
Table 5 shows the different APIs accessed by BotD and DataDome scripts on our honey site.
Table 5. Comparison of browser APIs read by DataDome and BotD
@@
Appendix C Combination of fingerprint attributes to evade DataDome
Algorithm 1 Algorithm to Detect Spatial Inconsistencies
1:  Input: Attribute categories $F$, Dataset containing requests $D$, Labels for requests $L$ (true if the request is from a bot, false for human)
2:  for all $f\in F$ do
3:     for all attribute pairs $\{f_{a},f_{b}\}\subseteq f$ do
4:        Filter $D$ where $L$ is false, creating $D^{\prime}$
6:        Sort the tuples in increasing order of $nv_{f_{b}}$
7:        for all $(v_{f_{a}},nv_{f_{b}})$ in the sorted order do
8:           if the combination is inconsistent then
9:              Label all rows in $D$ containing $(v_{f_{a}},v_{f_{b}})$ as true
10:           end if
11:        end for
12:     end for
13:  end for
Appendix D Algorithm to identify spatial inconsistencies
Algorithm 1 describes our algorithm to identify spatial inconsistencies.
Table 6. Inconsistencies Identified
Attribute Group Attributes Examples
@@ Appendix D Algorithm to identify spatial inconsistencies
Appendix E Inconsistencies Identified
Table 6 lists some examples of the inconsistencies that we identified for each attribute group in Table 7.
Table 7. Attribute Categories
Category Attributes
@@ Appendix E Inconsistencies Identified
Appendix F Attribute Categories for Inconsistency analysis
Table 7 list different categories of attributes used for inconsistency analysis.
Appendix G DataDome and BotD on Brave and Tor traffic
Roughly after the first 10 requests on each device, DataDome starts detecting all requests from Brave as bots resulting in a false positive rate of 41% on the 300 requests described in Section 7.5. BotD on the other hand does not detect any requests as bots.
Similar to FP-Inconsistent, DataDome detects all requests from Tor browser as bots while BotD does not detect any requests as bots. This further sheds light on the difficulty in distinguishing between Tor and bot traffic.